hi i need help seriously
\
g(x)=-2x+3 find the inverse of the function
log^381=x
5^x-4=625
27(3^x)=3^-1

Answer by stanbon(48564)

1) To form inverse of y=-2x+3
Interchange "x" and "y" to get:
x=-2y+3
Solve for "y", as follows:
y=(-x+3)/2
OR
y= -(1/2)x+(3/2)
Either of the last two is the inverse.
2) x=log^381
This problem does not make any sense
3)5^x-4=625
5^x=629
Take the log of both sides to get:
x(log 5)=log629
x=[log629]/[log5]
x= 4.004
4) 27(3^x)=3^(-1)
3^3)(3^x)=3^(-1)
3^(3+x)=3^(-1)
3+x=-1
x=-4
Cheers,
Stan H.